When 2 pipes are open, a tank is filled in 12 minutes. On one occasion, the first pipe was open for 10 minutes. Then the second pipe was opened and together they finished in 6 minutes. How long would it take each pipe by itself to fill the tank? Heeeeelp.
Answer by lwsshak3(6461)

When 2 pipes are open, a tank is filled in 12 minutes. On one occasion, the first pipe was open for 10 minutes. Then the second pipe was opened and together they finished in 6 minutes. How long would it take each pipe by itself to fill the tank
..
let x=minutes first pipe would take to fill the tank alone.
1/x=minute rate of first pipe
let y=minutes second pipe would take to fill the tank alone.
1/y=minute rate of second pipe
1/12=given minute rate when working together
sum of individual rates = rate when working together
first condition,
1/x+1/y=1/12
1/y=1/12-1/x
..
second condition
first pipe worked for 16 minutes with second pipe to finish the job
second pipe worked for 6 minutes with first pipe to finish the job
(1/x)*16+(1/y)*6=100%=1
(1/x)*16+(1/12-1/x)*6=1
16/x+6/12-6/x=1
10/x+1/2=1
10/x=1/2
x/10=2
x=20 minutes
1/y=1/12-1/x
1/y=1/12-1/20
LCD=60
60/y=5-3=2
y=30 minutes
ans:
It would take the first pipe 20 minutes to fill the tank by itself.
It would take the second pipe 30 minutes to fill the tank by itself.